# Decoupling Pulse Tube Vibrations from a Dry Dilution Refrigerator at milli-Kelvin Temperatures

### v1

@@ 3.2 Vibration Measurements
Using the vibration measurement setup described in section 2.1, we quantified the performance of the vibration decoupling system after the VDS comissioning run by comparing vibration spectra measured on the detector box with those taken on the cryostat’s mixing chamber. The corresponding acceleration spectra are shown in figure 5. The time domain data shows vibration levels obtained on the detector box which are about an order of magnitude lower than on the mixing chamber stage, when the pulse tube is running. The vibration level on the detector box with operating pulse tube is even lower than the one on the mixing chamber without pulse tube operation, demonstrating a vibration level during continuous cryostat operation smaller than what is typically achieved in an experimental volume only subject to environmental vibrations.
@@ 3.3 Cryogenic Detector Operation
As the NUCLEUS experiment aims to detect rare events from reactor anti-neutrino CENS interactions at recoil energies mainly below 100 eV, high stability together with ultra-low recoil energy thresholds are indispensable. This requires optimum detector performance with operating pulse tube at a stable base temperature below 10 mK. Using the presented vibration decoupling system, cryogenic detectors can be operated continuously below 10 eV baseline resolution. This value matches state-of-the-art performance of similar TES based calo-rimeters in wet cryostats [rothe2021low], while stable operation of this detector type directly on the mixing chamber of a DDR is not possible due to the high vibration levels. To illustrate performance and stability, we exemplarily present results from two different data taking runs during the commissioning phase of the NUCLEUS experiment. In the scope of an XRF calibration measurement in early 2023 a stable detector operation with running pulse tube over 20 days was obtained, as shown in figure 6.
@@ 4 Conclusion and Outlook
In this paper we have described the development of a vibration decoupling system for dry dilution refrigerators in the scope of the NUCLEUS experiment, with the aim of achieving continuous cryogenic detector operation undisturbed by both environmental and pulse tube related vibrations. It is based on the two main approaches of mounting the experimental volume on a frame of reference independent of the cryostat and decoupling remaining vibrations using a 1.8 m long spring pendulum. The resulting direct connection between room temperature and millikelvin temperatures poses a demanding thermal challenge, overcome by using a fine-tuned composition of radiation shields and thermal insulators. By coupling the vibration decoupling system only to selected stages of the cryostat with flexible thermal links, re-introduction of cryostat related vibrations into the system is minimized. Further studies in lowering the cross section of said thermal links by increasing the thermal coupling strength, and thus improving their flexibility, have the potential of further improving the achieved vibration levels and/or temperature stability.
Stable operating temperatures below 10 mK could be achieved in experiment repeatedly after less than two days of cooling time. Remaining vibration levels were demonstrated to be below what is typically achieved in an experimental volume only subject to environmental vibrations. It was demonstrated that the created cryogenic low vibration environment allows indefinite stable operation of TES based calorimeters at peak performance, uninfluenced by pulse tube vibrations. As such the developed vibration decoupling system facilitates the transition to dry dilution refrigerators for the cryogenic detector community without compromising performance due to pulse tube-induced vibrations.

### v2

@@ 3.2 Vibration Measurements
Using the vibration measurement setup described in section 2.1, we quantified the performance of the vibration decoupling system after the VDS comissioning run by comparing vibration spectra measured on the detector box with those taken on the cryostat’s mixing chamber. The corresponding acceleration spectra are shown in figure 5. The time domain data shows vibration levels obtained on the detector box which are about an order of magnitude lower than on the mixing chamber stage, when the pulse tube is running. The vibration level on the detector box with operating pulse tube is even lower than the one on the mixing chamber without pulse tube operation. This demonstrates a vibration level obtained during continuous cryostat operation below typical environmental vibrations.
@@ 3.3 Cryogenic Detector Operation
As the NUCLEUS experiment aims to detect rare events from reactor anti-neutrino CENS interactions at recoil energies mainly below 100 eV, high stability together with ultra-low recoil energy thresholds are indispensable. This requires optimum detector performance with operating pulse tube at a stable base temperature below 10 mK. Using the vibration decoupling system presented, cryogenic detectors can be operated continuously below 10 eV baseline resolution, while the operation of the same detectors directly on the mixing chamber was not possible due to the high vibration level caused by the pulse tube. This value matches state-of-the-art performance of similar TES based calorimeters in wet cryostats [rothe2021low]. To illustrate performance and stability, we exemplarily present results from two different data taking runs during the commissioning phase of the NUCLEUS experiment. In early 2023 an XRF calibration measurement was conducted [XRF]. The setup consisted of a 55Fe source that irradiated a two-stage target arrangement. This causes emission of calibration lines from 677 eV to 6.5 keV. The source was used to calibrate a 0.75 g CaWO4 crystal. With the entire assembly mounted on the vibration decoupling system presented in this work, a stable detector operation with running pulse tube was obtained over 20 days, as shown in figure 6.
@@ 4 Conclusion and Outlook
In this paper we have described the development of a vibration decoupling system for dry dilution refrigerators in the scope of the NUCLEUS experiment, with the aim of achieving continuous cryogenic detector operation undisturbed by both environmental and pulse tube related vibrations. It is based on the two main approaches of mounting the experimental volume on a frame of reference independent of the cryostat and decoupling remaining vibrations using a 1.8 m long spring pendulum. The resulting direct connection between room temperature and millikelvin temperatures poses a demanding thermal challenge, overcome by using a fine-tuned composition of radiation shields and thermal insulators. By coupling the vibration decoupling system only to selected stages of the cryostat with flexible thermal links, reintroduction of cryostat related vibrations into the system is minimized. Further studies in lowering the cross section of said thermal links by increasing the thermal coupling strength, and thus improving their flexibility, have the potential of further improving the achieved vibration levels and/or temperature stability.
Stable operating temperatures below 10 mK could be achieved in experiment repeatedly after less than two days of cooling time. The remaining vibration levels were shown to be below what is typically achieved in an experimental volume subject to environmental vibrations only. It was demonstrated that the created cryogenic low vibration environment allows indefinite stable operation of TES-based gram-scale calorimeters at peak performance, uninfluenced by pulse tube vibrations. The presented approach can also be applied to detectors with a higher mass. Although an overall larger mass would further improve axial decoupling (see equation 2.3), for the case of cryogenic calorimeters an increased mass leads to slower pulse time constants [probst1995model]. This can cause a shift of the detector ROI to lower frequencies. Given the measured VDS attenuation above 10 Hz in axial and 5 Hz in radial directions, the presented setup is expected to perform comparably in combination with detectors featuring moderately increased pulse time constants. For extremely slow pulses, additional optimization of the thermal links in order to further reduce low-frequency vibrations may be required. As such, the developed vibration decoupling system facilitates the transition to dry dilution refrigerators for the cryogenic detector community without compromising performance due to pulse tube-induced vibrations.